OpenStudy (anonymous):

Help solve!!! the quantity of x to the four thirds power, over x to the five sixths power the sixteenth root of the quantity of x times x to the third times x to the fourth

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@zimmah

OpenStudy (ranga):

Is this the first problem? \[|\large \frac{ x^{4/3} }{ x^{5/6} }\]

OpenStudy (anonymous):

yes

OpenStudy (ranga):

Use the identity: \[\large \frac{ x^m }{ x^n} = x^{m-n}\]

OpenStudy (anonymous):

idk how

OpenStudy (ranga):

\[\large \frac{ x^{4/3} }{ x^{5/6} } = x^{4/3 - 5/6} = x^{8/6 - 5/6} = x^{3/6} = x^{1/2}\]

OpenStudy (anonymous):

the second one is 16 on the outside of the radical symbol and x times x^3 times x^4 under the radical symbol

OpenStudy (jdoe0001):

morganfaith__ keep in mind that -> \(\bf \Large{ a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}} \qquad \qquad a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\\ \quad \\ \quad \\ \cfrac{x^{\frac{4}{3}}}{x^{\frac{5}{6}}}\implies \cfrac{x^{\frac{4}{3}}}{1}\cdot \cfrac{1}{x^{\frac{5}{6}}}\implies \cfrac{x^{\frac{4}{3}}}{1}\cdot x^{-\frac{5}{6}}\\ \quad \\ \implies x^{\frac{4}{3}+(-\frac{5}{6})}\implies x^{\frac{4}{3}-\frac{5}{6}}}\)

OpenStudy (anonymous):

what about the second one

OpenStudy (anonymous):

yes

OpenStudy (ranga):

x^3 times x^4 are both inside the radical and a small 16 is to the left of the radical?

OpenStudy (ranga):

\[\Large \sqrt[16]{x^3 \times x^4}?\]

OpenStudy (anonymous):

but its x*x^3*x^4 under the radical symbol

OpenStudy (jdoe0001):

\(\huge \bf \sqrt[16]{x\cdot x^3\cdot x^4}\quad ?\)

OpenStudy (anonymous):

jdoe has it right

OpenStudy (jdoe0001):

well, keep in mind that \(\Large \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\)

OpenStudy (jdoe0001):

change all "radicand" to rational exponentials, and add away

OpenStudy (anonymous):

to what?

OpenStudy (jdoe0001):

or .. ahemm... well. just add inside the radicand, and then change to rational exponent

OpenStudy (anonymous):

so it would be 3x^7?

OpenStudy (anonymous):

of just x^7

OpenStudy (jdoe0001):

\(\bf \large \sqrt[16]{x\cdot x^3\cdot x^4}\implies \sqrt[16]{x^1\cdot x^3\cdot x^4}\)

OpenStudy (ranga):

x * x^3 * x^4 = x^(1+3+4) = ?

OpenStudy (anonymous):

8?

OpenStudy (ranga):

x^8 Taking the 16th root is same as raising to the exponent (1/16)

OpenStudy (anonymous):

so the answer is x^8

OpenStudy (ranga):

No. (x^8)^(1/16) = ?

OpenStudy (jdoe0001):

\(\Large \bf \sqrt[{\color{red}{ 16}}]{x\cdot x^3\cdot x^4}\implies \sqrt[{\color{red}{ 16}}]{x^1\cdot x^3\cdot x^4}\implies \sqrt[{\color{red}{ 16}}]{x^\square }\implies x^{\frac{\square }{{\color{red}{ 16}}}}\)

OpenStudy (anonymous):

so x^8/16

OpenStudy (ranga):

simplify the exponent.

OpenStudy (anonymous):

which is x^1/2?

OpenStudy (ranga):

Yes. The answers to both problem is x^(1/2) or square root of x.

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (ranga):

You are welcome.